OpenStudy (anonymous):

Simplify. (If it is already in simplest form, enter the given expression.) x^2 − 16/ a(x + 4) − b(x + 4) so far I have = (x+4)(x-4)/a(x+4)-b(x+4) is there more to simplify this answer?

OpenStudy (anonymous):

\[\frac{ x^2-16 }{ a(x+4)-b(x+4) } =\frac{ (x+4)(x-4) }{ a(x+4)-b(x+4) }\]

OpenStudy (anonymous):

thought the answer would be (x+4)(x-4)/a-b (but you cannot simplify the denominator since it has the minus/subtraction symbol correct? if you could help me understand the rule greatly appreciate it )

OpenStudy (anonymous):

Can anybody see this?